OpenStudy (anonymous):

The number of milligrams of a drug that remains in a patient's system after t hours is given by the function A(t) = Iert. Jordan was given 350 milligrams of medicine which leaves his bloodstream at a rate of 20%. How much of the medicine remains in his system after 4 hours?

OpenStudy (kropot72):

Is this the function?\[A(t)=Ie ^{-rt}\]

OpenStudy (anonymous):

yea

OpenStudy (kropot72):

\[A(4\min)=350\times e ^{-(0.2\times 4)}=\frac{350}{e ^{0.8}}\] Can you calculate the answer?

OpenStudy (anonymous):

actually i cant... can you walk me through it?

OpenStudy (kropot72):

You have to divide 350 by the exponential number 'e' raised to the power of 0.8. Can you find the value of the exponential number 'e' raised to the power of 0.8? A suitable calculator will do it

OpenStudy (anonymous):

what is the value of e?

OpenStudy (anonymous):

what is the value of e?

OpenStudy (anonymous):

or is that unknown?

OpenStudy (kropot72):

An approximate value for 'e' is 2.7183

OpenStudy (anonymous):

im getting some wierd answers that dont make any sense

OpenStudy (anonymous):

im getting some wierd answers that dont make any sense

OpenStudy (anonymous):

im getting some wierd answers that dont make any sense

OpenStudy (kropot72):

\[e ^{0.8}=2.2255\] Amount of medicine in bloodstream after 4 hours = 350/2.2255 milligrams